OpenStudy (anonymous):

(3a^3 − 5b^3) + _?_ a^3 +_?_ b^3 = (a^3+ b^3).

Mehek (mehek14):

\(3a^3-a^3=?\)

Mehek (mehek14):

i's basically like subtracting 3-1 but you keep \(a^3\)

TheSmartOne (thesmartone):

Hint: \(\sf a^3 + b^3 = (a+b) (a^2-ab+b^2)\)